OpenStudy (ttop0816):

In triangle ABC, a = 8, b = 6, and C = 60°. Find B. 14° 46° 74° 106° <

OpenStudy (yttrium):



OpenStudy (yttrium):

you may solve c first by applying cosine law: c = sqrt( 8^2 + 6^2 - cos60)

OpenStudy (yttrium):

After solving c, you may now solve for angle B using sine law: sin60/c = sinB/6 Hence, B = arcsin[(6sin60)/c]